OpenStudy (anonymous):

A sector of a circle with a 120 degree arc has an area of 75(pi)in^2. Find the radius.

OpenStudy (kira_yamato):

θ = 120˚ = π/3 rad A = rθ 75 = r/3 r = 225 in

OpenStudy (anonymous):

is the radius 225?

OpenStudy (kira_yamato):

I guess...

OpenStudy (king):

isnt it r^2=225 therefore,r=15 in ??

OpenStudy (anonymous):

did u square root 225?

OpenStudy (kira_yamato):

Sorry... Use wrong formula... I rework it...

OpenStudy (king):

yeah..

OpenStudy (anonymous):

u dnt have to rework.. it was correct u only had to do the square root of 225 at the end. thanks :)

OpenStudy (kira_yamato):

A = 0.5 r^2 θ 75π = 0.5 (π/3)(r^2) 75 = (r^2)/6 r^2 = 450 r = √450 = 15√2